The sum of the digits of a two digit number is 7. When the digits are reversed the value of the number increased by 27. Find the number

Found 2 solutions by fractalier, solver91311:
Answer by fractalier(6550)

Call the tens digit x and the ones digit y.
Thus x + y = 7
Originally the value is 10x + y.
Reversed its value is 10y + x. This is 27 more than the original, so
10y + x = 10x + y + 27
or
-9x + 9y = 27
or
-x + y = 3 now add the first equation and get
x + y = 7
2y = 10
y = 5
x = 2
The original number was 25.

Answer by solver91311(24713)



Shortcut Method

If you reverse the digits of a two-digit number, the difference between the original two digit number and the new two-digit number is ALWAYS a multiple of 9. And the difference between the two digits is the value of the multiple. Hence, for your problem, where is the original 10s digit and is the original 1s digit:



And, since the original two-digit number must be smaller than the new two-digit number, meaning

.

Long Method

Same definitions of and

Original two digit number:

New two digit number:







And



Either way, solve the 2X2 system for and

John

My calculator said it, I believe it, that settles it